OpenStudy (anonymous):

A 20kg cart is pushed with a force of 175N at an angle of 25degrees. if it starts from rest, how far does it travel in 7 seconds?

OpenStudy (anonymous):

if i am true ... Equation of conservation \[Ec _{A}+W(F)=EC _{B}\] \[1/2mv _{A}^{2}+Fdcos25=1/2mv _{B}^{2}\] \[Ec _{A}=0\] \[V _{B}^{2}=(2m)(Fdcos25)\] and continue ......

OpenStudy (anonymous):

the effective force in the direction of motion is fcos25. we know that f=ma; a=f/m; d=ut + a/2*t^2 as initial velocity is zero , u is 0. calculate d from the above equation.